# Anosov representations, strongly convex cocompact groups and weak eigenvalue gaps

### v3

@@ 10. Examples and counterexamples
Necessity of regularity of $\partial\Omega$ in Proposition 8.1. Let $n\geqslant 2$ and $\Gamma$ be a convex cocompact subgroup of $\mathsf{SU}(n,1)\subset\mathsf{SL}_{n+1}(\mathbb{C})$. Let $\tau_{2}:\mathsf{SL}_{n+1}(\mathbb{C})\xhookrightarrow{}\mathsf{SL}_{2n+2}(\mathbb{R})$ be the standard inclusion defined as
$\tau_{2}(g)=\begin{pmatrix}\textup{Re}(g)&-\textup{Im}(g)\\ \textup{Im}(g)&\textup{Re}(g)\end{pmatrix},\ \ g\in\mathsf{SL}_{n+1}(\mathbb{C}).$
The group $\textup{sym}^{2}\big{(}\tau_{2}(\Gamma)\big{)}\subset\mathsf{SL}_{2n+2}(\mathbb{R})$ is a $P_{2}$-Anosov subgroup and there exist $J,k>0$ such that
$J^{-1}(\gamma\cdot\delta)_{e}-k\leq\big{(}\textup{sym}^{2}(\tau_{2}(\gamma))\cdot\textup{sym}^{2}(\tau_{2}(\delta))\big{)}_{\varepsilon_{1}}\leq J(\gamma\cdot\delta)_{e}+k$
for every $\gamma,\delta\in\Gamma$. Moreover, $\textup{sym}^{2}\big{(}\tau_{2}(\Gamma)\big{)}$ preserves a properly convex domain in $\mathbb{P}\big{(}\textup{Sym}^{2}\mathbb{R}^{2n+2}\big{)}$ but it cannot preserve a strictly convex domain since it is not $P_{1}$-divergent. Similar counterexamples are given by convex cocompact subgroups of the rank 1 Lie group $\mathsf{Sp}(n,1)\subset\mathsf{GL}_{n+1}(\mathbb{H})$.
@@
Necessity of transversality in Theorem 1.1 in the Zariski dense case. There exists a Zariski dense representation $\rho_{1}:\pi_{1}(S_{g})\rightarrow\mathsf{PSL}_{4}(\mathbb{R})$ which admits a pair of continuous $\rho_{1}$-equivariant maps $\xi^{+}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{4})$ and $\xi^{-}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathsf{Gr}_{3}(\mathbb{R}^{4})$ but is not $P_{1}$-Anosov.
Let $M^{3}$ be a closed hyperbolic $3$-manifold fibering over the circle (with fiber $S_{g}$) which also contains a totally geodesic surface. By Johnson–Millson [25] the natural inclusion $j:\pi_{1}(M^{3})\xhookrightarrow{}\mathsf{PO}(3,1)$ admits a non-trivial Zariski dense deformation $j^{\prime}:\pi_{1}(M^{3})\rightarrow\mathsf{PSL}_{4}(\mathbb{R})$ which by can be chosen to be $P_{1}$-Anosov thanks to the openess of Anosov representations (see [33, 23]). Let $\xi_{1}^{+}$ and $\xi_{1}^{-}$ be the Anosov limit maps of $j^{\prime}$ into $\mathbb{P}(\mathbb{R}^{4})$ and $\mathsf{Gr}_{3}(\mathbb{R}^{4})$ respectively. By the theorem of Cannon-Thurston [13] there exists a continuous, $\pi_{1}(S)$-equivariant map $\theta:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\partial_{\infty}\pi_{1}(M^{3})$. The restriction $\rho_{1}:=j^{\prime}|_{\pi_{1}(S_{g})}$ is Zariski dense, not a quasi-isometric embedding and $\xi_{1}^{+}\circ\theta$ and $\xi_{1}^{-}\circ\theta$ are continuous, non-transverse and $\rho_{1}$-equivariant maps. In addition, by [10], every finitely generated free subgroup $F$ of $\pi_{1}(S_{g})$ is a quasiconvex subgroup of $\pi_{1}(M^{3})$. Hence, $\iota^{\prime}|_{F}$ is $P_{1}$-Anosov and $\xi^{+}\circ\iota_{F}$ and $\xi^{-}\circ\iota_{F}$ are transverse.

### v4

@@ 10. Examples and counterexamples
Necessity of regularity of $\partial\Omega$ in Proposition 8.1. Let $n\geqslant 2$ and $\Gamma$ be a convex cocompact subgroup of $\mathsf{SU}(n,1)\subset\mathsf{SL}_{n+1}(\mathbb{C})$. Let $\tau_{2}:\mathsf{GL}_{n+1}(\mathbb{C})\xhookrightarrow{}\mathsf{GL}_{2n+2}(\mathbb{R})$ be the standard inclusion defined as
$\tau_{2}(h):=\begin{pmatrix}\textup{Re}(h)&-\textup{Im}(h)\\ \textup{Im}(h)&\textup{Re}(h)\end{pmatrix},\ \ h\in\mathsf{GL}_{n+1}(\mathbb{C}).$
Note that for every $h\in\mathsf{SU}(n,1)$, since $\sigma_{1}(\tau_{2}(h))=\sigma_{2}(\tau_{2}(h))=\sigma_{1}(h)$ and $\sigma_{i}(h)=1$ for $i=3,\ldots,2n$, the subgroup $\tau_{2}(\Gamma)\subset\mathsf{SL}_{2n+2}(\mathbb{R})$ is $P_{2}$-Anosov but not $P_{1}$-Anosov (in particular not $P_{1}$-divergent). In addition, since $\sigma_{1}(\textup{sym}^{2}(\tau_{2}(\gamma)))=\sigma_{1}(\tau_{2}(\gamma))^{2}=\sigma_{1}(\gamma)^{2}$ for every $\gamma\in\Gamma\subset\mathsf{SU}(n,1)$, we conclude that there exist $J,k>0$ such that
$J^{-1}(\gamma_{1}\cdot\gamma_{2})_{e}-k\leq\Big(\textup{sym}^{2}(\tau_{2}(\gamma_{1}))\cdot\textup{sym}^{2}(\tau_{2}(\gamma_{2}))\Big)_{\varepsilon_{1}}\leq J(\gamma_{1}\cdot\gamma_{2})_{e}+k$
for every $\gamma_{1},\gamma_{2}\in\Gamma$. Moreover, $\textup{sym}^{2}\big(\tau_{2}(\Gamma)\big)$ preserves a properly convex domain in $\mathbb{P}\big(\textup{Sym}^{2}\mathbb{R}^{2n+2}\big)$ but it cannot preserve a strictly convex domain since $\textup{sym}^{2}(\tau_{2}(\Gamma))\subset\mathsf{SL}(\textup{Sym}^{2}(\mathbb{R}^{d}))$ is not $P_{1}$-divergent.
Similar examples are given by convex cocompact subgroups of the rank 1 Lie group $\mathsf{Sp}(n,1)\subset\mathsf{GL}_{n+1}(\mathbb{H})$, where $\mathbb{H}=\mathbb{C}\oplus\mathbb{C}j$ are Hamilton’s quaternions. By using the standard embedding $\tau_{4}:\mathsf{GL}_{n+1}(\mathbb{H})\xhookrightarrow{}\mathsf{GL}_{4n+4}(\mathbb{R})$, $\tau_{4}(C+Dj)=\tau_{2}\Big(\begin{pmatrix}[0.3]C&-D\\ \overline{D}&\overline{C}\end{pmatrix}\Big)$, where $C+Dj\in\mathsf{GL}_{n+1}(\mathbb{H})$, $C,D\in\mathsf{Mat}_{n+1}(\mathbb{C})$, for any convex cocompact subgroup $\Delta\subset\mathsf{Sp}(n,1)$, $\tau_{4}|_{\Delta}$ is $P_{4}$-Anosov but not $P_{1}$-Anosov. In addition, there are $R,r>1$ such that for every $h_{1},h_{2}\in\Delta$,
$R^{-1}(h_{1}\cdot h_{2})_{e}-r\leq\Big(\textup{sym}^{2}(\tau_{4}(h_{1}))\cdot\textup{sym}^{2}(\tau_{4}(h_{2}))\Big)_{\varepsilon_{1}}\leq R(h_{1}\cdot h_{2})_{e}+r.$
@@
Necessity of transversality in Theorem 1.1 in the Zariski dense case. There exists a Zariski dense representation $\rho_{1}:\pi_{1}(S_{g})\rightarrow\mathsf{PSL}_{4}(\mathbb{R})$ which is not $P_{1}$-Anosov but it admits a pair of continuous $\rho_{1}$-equivariant maps $\xi^{+}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{4})$ and $\xi^{-}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathsf{Gr}_{3}(\mathbb{R}^{4})$.
Let $M^{3}$ be a closed hyperbolic $3$-manifold which contains a totally geodesic surface. By [2], up to replacing $M^{3}$ with a finite cover, we may also assume that $M^{3}$ fibers over the circle (with fiber $S_{g}$). By [27] the natural inclusion $j:\pi_{1}(M^{3})\xhookrightarrow{}\mathsf{PO}(3,1)$ admits a non-trivial Zariski dense deformation $j^{\prime}:\pi_{1}(M^{3})\rightarrow\mathsf{PSL}_{4}(\mathbb{R})$ which can be chosen to be $P_{1}$-Anosov, thanks to the openess of Anosov representations (see [36, 25]). Let $\xi_{1}^{+}$ and $\xi_{1}^{-}$ be the Anosov limit maps of $j^{\prime}$ into $\mathbb{P}(\mathbb{R}^{4})$ and $\mathsf{Gr}_{3}(\mathbb{R}^{4})$ respectively. By the theorem of Cannon-Thurston [15] there exists a continuous, $\pi_{1}(S_{g})$-equivariant map $\theta:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\partial_{\infty}\pi_{1}(M^{3})$. The restriction $\rho_{1}:=j^{\prime}|_{\pi_{1}(S_{g})}$ is Zariski dense, not a quasi-isometric embedding and $\xi_{1}^{+}\circ\theta$ and $\xi_{1}^{-}\circ\theta$ are continuous, non-transverse and $\rho_{1}$-equivariant maps. In addition, by [12], every finitely generated free subgroup $F$ of $\pi_{1}(S_{g})$ is a quasiconvex subgroup of $\pi_{1}(M^{3})$. Hence, $\iota^{\prime}|_{F}$ is $P_{1}$-Anosov and $\xi^{+}\circ\iota_{F}$ and $\xi^{-}\circ\iota_{F}$ are transverse.
By [37, Thm. 7.5], there are also examples of Zariski dense representations $\psi:\mathsf{\Delta}\rightarrow\mathsf{SL}_{3}(\mathbb{R})$ of triangle reflection groups $\mathsf{\Delta}$, which admit continuous, $\psi$-equivariant, injective maps $\xi^{1}:\partial_{\infty}\mathsf{\Delta}\rightarrow\mathbb{P}(\mathbb{R}^{3})$, $\xi^{2}:\partial_{\infty}\mathsf{\Delta}\rightarrow\mathsf{Gr}_{2}(\mathbb{R}^{3})$ (hence $\psi$ is discrete and faithful), but $\psi$ is not $P_{1}$-Anosov.